# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ 4.3.3 Low Permanent-degree Reductions
We further introduce two reduction rules to reduce high-degree vertices that have only one or two permanent neighbors.
@@
If there exists a vertex $u$ that has exactly one permanent neighbor $x$ and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(u,v))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, and $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
@@
Let $u$ be a vertex with exactly two permanent neighbors $x,y$, that $p(x,y)=M$, and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(u,v))$. W.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$.
Case 1: If $w(u)\geq\tilde{w}(x)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,x,y\}$, and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
Case 2: If $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,y\}$, $w(x)\coloneqq w(x)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, and $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
Case 3: If $w(u)<\tilde{w}(y)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$, and $p(y,v)\coloneqq p(y,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(y)$.
@@ 4.4 Vertex-pair Reductions
If there is a permanent edge $(u,v)$ such that $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\min(w^{+}(N_{r}(u)\setminus N_{p}(v)),\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x)))$, then $\alpha(G)=\alpha(G\setminus\{v\})$.
@@
If there is a permanent edge $(u,v)$ such that $w(v)\geq w^{+}(N(v))-\max(0,w(u))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus(N_{p}(u)\cap N_{p}(v))$.
@@
If there exist two vertices $u,v$ such that $N_{p}(u)=N_{p}(v)$, $w(u)\geq\sum_{x\in N_{r}(u)}\max(0,p(u,x))$ and $w(v)\geq\sum_{x\in N_{r}(v)}\max(0,p(v,x))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}$ is obtained by folding $u$ and $v$ into a vertex $v^{\prime}$, $w(v^{\prime})\coloneqq w(u)+w(v)-p(u,v)$, $p(u,x)\coloneqq M$ for each $x\in N_{p}(u)$, and $p(v^{\prime},x)\coloneqq p(u,x)+p(v,x)$ for each $x\in N_{r}(v)\cup N_{r}(u)$.
@@ 5 Reduction-driven Local Search
$S_{init}\leftarrow\varnothing$; /* initial an empty solution */
$G,S_{init}\leftarrow\textit{Complete\_Reduction}(G,S_{init})$; /* Pre-process to obtain a kernel graph via $R1$-$R14$/
@@
       /* construct an initial solution using $R1$-$R14$/
@@ 3       $S\leftarrow\textit{Neighborhood\_Search\_with\_Partial\_Reduction}(G,S)$;
The RLS begins by pre-processing the original graph $G$. It uses the Complete_Reduction procedure which applies all the reduction rules to obtain a kernel graph. Due to the correctness of the reduction rules, solving the GIS problem in the original graph is equivalent to solving it in this kernel graph. The $S_{init}$ is a partial vertex set that must be in the optimal solution.
The remaining steps of RLS follow the traditional iterated local search algorithm (Glover and Laguna (1998), Lourenço et al. (2019)): At each iteration, it supplements $S_{init}$ with a random generalized independent set obtained by Random_Peeling, then uses Neighborhood_Search_with_Partial_Reduction to improve the current solution (line 1)-1) . The current solution is denoted by $S$ and the best-known solution visited by the iterative search is kept in $S^{*}$ (line 1). The iteration continues until the cutoff time limit is met. Compared with traditional local search, RLS seamlessly integrates reduction techniques into the pre-processing, the initial solution building, and the neighborhood searching procedures.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S\subseteq V$
@@
8                   if the condition of $Ri$ is satisfied by checking $elem$ then
9                         change the structure, profits and penalties of $G$ and set $S$ according to $Ri$;
10                         update the Candj if $Rj$ reachable from $Ri$ in Figure 4;
@@
12                  Remove $elem$ from Candi;
@@ 5.1 Complete Reduction as Pre-processing
The Complete_Reduction algorithm exhaustively applies the 14 reduction rules in Section 4. A naive implementation is to check every rule independently, process the graph when applicable, and stop when no rules apply. The times of checking all these rules is $O(14(|V|+|E|))$ as it requires scanning the entire graph for each rule. Given the large size of real-world graphs, this is inefficient in practice. We propose a more efficient implementation which incrementally detects applicable rules. Specifically, we observe that rules are interdependent, with one application triggering others. For example, if only the condition of $R1$ holds in graph $G$ (i.e., there exists an edge $(u,v)$ with $p(u,v)=0$), the edge $(u,v)$ is removed, making R3-15 potentially applicable but not the other rules. Hence, we make an in-depth analysis of this dependency relationship for all these rules, and we show the result in Figure 4.
In Algorithm 2, we implement this reduction algorithm. It takes a graph $G$ as its input and outputs a kernel graph and a set $S\subseteq V$. In detail, it initializes a special candidate set Candi which contains either vertices, edges, or vertex pairs, for a rule $Ri$ where $i=1,...,14$. Lines 1-4 show how Candi is initialized. When an element $elem\in\text{Cand}_{i}$ satisfies the condition of $Ri$, the current graph $G$ is changed and the current set $S$ is updated according to $Ri$. Due to the change of graph $G$, if another reduction rule $Rj$ can be applied (i.e. $Rj$ is reachable from $Ri$ in the dependency graph in Figure 4), then the vertices, edges, or vertex pairs which are “neighbor elements” of $elem$ are moved into Candj. Consequently, the algorithm checks whether each reduction rule is satisfied by tracking the candidate set in an incremental manner.
@@
5.2 Random Peeling for Initial Solutions
Random-peeling is a simple random procedure for generating an initial solution. It starts with an empty vertex set $S$. Then when $V\setminus N_{p}(S)$ is not empty (meaning that $S$ can possibly be a larger generalized independent set), a vertex $u$ is picked randomly from $V\setminus N_{p}(S)$ and $u$ is moved into $S$. Due to the removal of $u$ from $G$, the Complete_Reduction is used again to further shrink the remaining graph $G$ (line 3). When $V\setminus N_{p}(S)$ becomes empty, we further extend $S$ by randomly picking a vertex $v$ such that $nb(S\cup\{v\})>nb(S)$. The $S$ is returned when it becomes a maximal generalized independent set.
@@ 3       $u\leftarrow$ randomly select a vertex from $V\setminus N_{p}[S]$;
5      $G,S\leftarrow\textit{Complete\_Reduction}(G\setminus\{u\},S)$;
@@ 5.3 Neighborhood Search
To improve the incumbent solution, we use the iterated tabu search with three move operators.
Move operators. In RLS, there are three move operators that update the current solution: ADD, SWAP, and the novel REDUCE operator based on two reduction rules. Given the current solution $S\subseteq V$, the ADD$(v)$ operator displaces a vertex $v$ (if any) from $V\setminus N_{p}[S]$ to the $S$ and the SWAP$(u,v)$ operator adds a vertex $v\in N_{p}(S)$ if $|N_{p}(v)\cap S|=1$ and removes the only vertex $u\in N_{p}(v)\cap S$ from the current solution $S$. Indeed, ADD and SWAP are often used in existing local search methods such as Zhou and Hao (2017), Jin and Hao (2015). The REDUCE$(V\setminus N_{p}[S])$ operator applies $R3$ and $R4$ to include candidate vertices that are definitely in a set in $MGIS(G)$.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S$, search depth $L\in\mathbb{N}$, perturbation parameter $\epsilon\in(0,1)$
@@ 1 $S^{*}\leftarrow S$, $tabu\_list\leftarrow\varnothing$;
3       $M_{1}\leftarrow V\setminus N_{p}[S]$, $M_{2}\leftarrow\{v\in N_{p}(S)\mid u\in S,(u,v)\in E_{p},|N_{p}(v)\cap S|=1\}\setminus tabu\_list$
@@ 4       $v\leftarrow$ a vertex with the largest move gain from $M_{1}\cup M_{2}$;
       if $v\in M_{1}$ then  $S\leftarrow S\cup\{v\}$, $M_{1}\leftarrow M_{1}\setminus\{v\}$ /* add $v$ into $S$/
@@
             $u\leftarrow N_{p}(v)\cap S$, $S\leftarrow S\cup\{v\}\setminus\{u\}$; /* Swap $v$ and $u$/
6             add $u$ into $tabu\_list$ with tabu tenure $T_{u}=10+random(0,|M_{2}|)$
       $REDUCE(M_{1})$ and reassign $S$; /* apply $R3$ and $R4$/
@@
Fast Calculation of Move Gain. For each vertex $u\in V$, we maintain $B_{u}=w(u)-\sum_{v\in S,(u,v)\in E}p(u,v)$. Then, the move gain of $ADD(v)$, which is the gain of the net benefit after applying a $ADD(v)$, is $B_{v}$, and the move gain of $SWAP(u,v)$ is $B_{v}-B_{u}$. After an $ADD(v)$ move, for each $u\in N_{r}(v)$, we update $B_{u}$ as $B_{u}\coloneqq B_{u}-p(u,v)$ After a $SWAP(u,v)$ move, for each $x\in N_{r}(u)$, $B_{x}\coloneqq B_{x}+p(x,u)$ and for each $y\in N_{r}(v)$, $B_{y}\coloneqq B_{y}-p(y,v)$. The time complexity of performing $ADD$ or $SWAP$ is $O(\Delta)$, while the REDUCE operator operates exactly. The overall complexity of applying all three operators is $O(|E|)$. Notably, the REDUCE operator is not specific to the GIS problem and can enhance the efficiency of local search in other optimization problems.
Neighborhood Search with Tabu and Perturbation. The neighborhood search in RLS iteratively performs the three basic operators above to improve the current solution $S$, as shown in Algorithm 4.
At each iteration in lines 4-4, it first builds $M_{1}$, the set of vertices that can be added into $S$ by $ADD$, and $M_{2}$, the set of vertices that can be swapped into $S$ by $SWAP$. By evaluating $M_{1}$ and $M_{2}$, the $ADD$ or $SWAP$ with the largest move gain is applied, as presented in lines 4-8. After performing the ADD or SWAP operator, we additionally use the REDUCE operator to update $M_{1}$. This decreases the search space of the $ADD$ operator because $V\setminus N_{p}(S)$ is much larger than $S$ in practice.
To avoid short-term search cycles, a tabu list is used to exclude recently visited vertices for successive iterations (line 4). The tabu tenure for a vertex is set to $10+random(0,|M_{2}|)$, as commonly used in local search algorithms (Wu et al. 2023, Zhou and Hao 2017, Jin and Hao 2015). To further avoid local optima, a perturbation step is applied during neighborhood search (lines 4-4). Specifically, when we detect that no operators can be applied or the best solution remains unchanged after a certain number of iterations, we remove a portion of low-contribution vertices. The whole neighborhood search terminates when the consecutive number of iterations without improvement reaches the search depth parameter $L$ (line 4).

### v2

@@ 4.3.3 Low Permanent-degree Reductions
We further introduce two reduction rules to reduce high-degree vertices that are connected by only one or two vertices by permanent edges.
@@
If there exists a vertex $u$ such that $u$ has exactly one neighbor $x\in N_{p}(u)$ and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,v\}\mid v\in N_{r}(u)\setminus N_{p}(x)\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$, and $p^{\prime}(\{x,v\})\coloneqq p(\{x,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(x)$. $u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
Let $u$ be a vertex with exactly two neighbors $x,y\in N_{p}(u)$, $\{x,y\}\in E_{p}$, and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))$. W.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$.
Case 1: If $w(u)\geq\tilde{w}(x)$, then $u$ must be included in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,x,y\}]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
Case 2: If $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ is not in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u,y\},E(G[V\setminus\{u,y\}])\cup\{\{x,v\}\mid v\in N_{r}(u)\setminus N_{p}(x)\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$, and $p^{\prime}(\{x,v\})\coloneqq p(\{x,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(x)$. $u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
Case 3: If $w(u)<\tilde{w}(y)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,v\}\mid v\in N_{r}(u)\setminus N_{p}(x)\}\cup\{\{y,v\}\mid v\in N_{r}(u)\setminus N_{p}(y)\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-w(u)$, $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$, $p^{\prime}(\{x,v\})\coloneqq p(\{x,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(x)$, and $p^{\prime}(\{y,v\})\coloneqq p(\{y,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(y)$. $u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@ 4.4 Vertex-pair Reductions
If there is an edge $\{u,v\}$ such that $\{u,v\}\in E_{p}$ and $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\min(w^{+}(N_{r}(u)\setminus N_{p}(v)),\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p\{u,x\}))$, then $v$ is not in at least one $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{v\}]$.
@@
If there is an edge $\{u,v\}$ such that $\{u,v\}\in E_{p}$ and $w(v)\geq w^{+}(N(v))-\max(0,w(u))$, then $N_{p}(u)\cap N_{p}(v)$ is not a subset of at least one $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G[V\setminus(N_{p}(u)\cap N_{p}(v))]$.
@@
If there exist two vertices $u,v$ such that $N_{p}(u)=N_{p}(v)$, $w(u)\geq\sum_{x\in N_{r}(u)}\max(0,p(\{u,x\}))$ and $w(v)\geq\sum_{x\in N_{r}(v)}\max(0,p(\{v,x\}))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}$ is obtained by folding $u$ and $v$ into a vertex $v^{\prime}$, $w(v^{\prime})\coloneqq w(u)+w(v)-p(\{u,v\})$, add edge $\{v^{\prime},x\}$ as a permanent edge in $G^{\prime}$ for each $x\in N_{p}(u)$, and add edge $\{v^{\prime},x\}$ as a removable edge in $G^{\prime}$ with $p^{\prime}(\{v^{\prime},x\})\coloneqq p(\{u,x\})+p(\{v,x\})$ for each $x\in N_{r}(v)\cup N_{r}(u)$. Both $u$ and $v$ belong to at least one $I\in MGIS(G)$ iff $v^{\prime}\in MGIS(G^{\prime})$.
@@ 5 Reduction-driven Local Search
$G,S_{init}\leftarrow\textit{Complete\_Reduction}(G)$; /* Pre-process to obtain a kernel graph and partial solution via $R1$-$R14$ */
@@
 /* construct an initial solution using $R1$-$R14$ */
@@ 3 $S\leftarrow\textit{Neighborhood\_Search\_with\_Partial\_Reduction}(G,S)$;
The RLS begins by pre-processing the original graph $G$. It uses the Complete_Reduction procedure which applies all the reduction rules to obtain a kernel graph, that is, a graph where no further reduction rules can be applied. Due to the correctness of the reduction rules, solving the GIS problem in the original graph is equivalent to solving it in this kernel graph. The $S_{init}$ is a partial vertex set that must be a subset of at least one optimal solution.
The remaining steps of the RLS follow the traditional iterated local search algorithm (Glover and Laguna (1998), Lourenço et al. (2019)): At each iteration, it supplements $S_{init}$ with a random generalized independent set obtained by Random_Peeling, then uses Neighborhood_Search_with_Partial_Reduction to improve the current solution (line 1-1) . The current solution is denoted by $S$ and the best-known solution visited by the iterative search is kept in $S^{*}$ (line 1). The iteration continues until the cutoff time limit is met. Compared with traditional local search, the RLS seamlessly integrates reduction techniques into the pre-processing, the initial solution building, and the neighborhood searching procedures.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties
@@
8       if the condition of $Ri$ is satisfied by checking elemelem then
9            change the structure, profits and penalties of $G$ and set $S$ according to RiRi;
10            update the Candj if RjRj reachable from $Ri$ in Figure 4;
@@
12        Remove elemelem from Candi;
@@ 5.1 Complete Reduction as Pre-processing
The Complete_Reduction algorithm exhaustively applies the 14 reduction rules described in Section 4. A naive implementation is to check every rule independently, process the graph when applicable, and stop when no rules apply. The times for checking all these rules are bounded by $O(|V|+|E|)$ as it requires scanning the entire graph for each rule. Given the large size of real-world graphs, this is inefficient in practice. We propose a more efficient implementation that incrementally detects applicable rules. Specifically, we observe that rules are interdependent, with one application triggering others. For example, if only the condition of $R1$ is satisfied in graph $G$ (i.e., there exists an edge $\{u,v\}\in E_{r}$ with $p(\{u,v\})=0$), then the edge $\{u,v\}$ is removed, potentially enabling the application of R3R3-$14$, while the remaining rules remain inapplicable. Hence, we make an in-depth analysis of this dependency relationship for all these rules, and we show the result in Figure 4.
In Algorithm 2, we implement this reduction algorithm. It takes a graph $G$ as its input and outputs a kernel graph and a set $S\subseteq V$. In detail, it initializes a special candidate set Candi, which contains either vertices, edges, or vertex pairs, for a rule $Ri$ where i=1,,14i=1,...,14. Lines 1-4 show how Candi is initialized. When an element $elem\in\text{Cand}_{i}$ satisfies the condition of $Ri$, the current graph $G$ is changed and the current set $S$ is updated according to $Ri$. Along with the change of graph $G$, if another reduction rule $Rj$ can be applied (i.e., $Rj$ is reachable from $Ri$ in the dependency graph in Figure 4), then the vertices, edges, or vertex pairs which are “neighbor elements” of elemelem are moved into Candj. The process continues recursively until all candidate sets are empty. Therefore, the algorithm checks whether each reduction rule is satisfied by tracking the candidate set in an incremental manner.
In our implementations, $R1$ and $R2$ are prioritized because the removal of edges may enable more reduction rules. $R3$-$R6$ are assigned secondary priority, as they directly identify vertices that are in at least one set in MGIS(G), which may further shrink the candidate set. $R7$-$R11$ are of the third priority because they can identify vertices that are in at least one set in MGIS(G) in some cases. $R12$-$R14$ are put at the last because they are not able to identify such vertices.
@@ 3 $u\leftarrow$ randomly select a vertex from $V\setminus N_{p}[S]$;
5 $G,S\leftarrow\textit{Complete\_Reduction}(G[V\setminus\{u\}],S)$;
@@
5.2 Random Peeling for Initial Solutions
Random-peeling is a simple random procedure for generating an initial solution. It starts with an empty vertex set $S$. When $V\setminus N_{p}[S]$ is not empty (meaning that $S$ can possibly be a larger generalized independent set), a vertex $u$ is picked randomly from $V\setminus N_{p}[S]$ and $u$ is moved into $S$. After the removal of $u$ from $G$, the Complete_Reduction is used again to further shrink the remaining graph $G$ (line 3). When $V\setminus N_{p}[S]$ becomes empty, we further extend $S$ by randomly picking a vertex $v$ from the randomly removed vertices such that $v\notin N_{p}[S]$ and $nb(S\cup\{v\})>nb(S)$. The $S$ is returned when there is no other generalized independent set $S^{\prime}$ such that $S\subset S^{\prime}$ and $nb(S^{\prime})\geq nb(S)$. Clearly $S$ is a maximal generalized independent set.
@@ 5.3 Neighborhood Search
To improve the current solution, we use an iterated tabu search with three move operators.
[…]
Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S$, search tolerance $L\in\mathbb{N}$, perturbation parameter $\epsilon\in(0,1)$
@@ 1 $S^{*}\leftarrow S$, $tabu\_list\leftarrow\varnothing$;
3 $M_{1}\leftarrow V\setminus N_{p}[S]$, $M_{2}\leftarrow\{v\in N_{p}(S)\mid u\in S,\{u,v\}\in E_{p},|N_{p}(v)\cap S|=1\}\setminus tabu\_list$
@@ 4 $v\leftarrow$ a vertex with the largest move gain from $M_{1}\cup M_{2}$;
 if $v\in M_{1}$ then $S\leftarrow S\cup\{v\}$, $M_{1}\leftarrow M_{1}\setminus\{v\}$ /* add $v$ into $S$ */
@@
    $u\leftarrow N_{p}(v)\cap S$, $S\leftarrow S\cup\{v\}\setminus\{u\}$; /* Swap $v$ and $u$ */
6      add $u$ into tabu_listtabu\_list with tabu tenure $T_{u}=10+random(0,|M_{2}|)$
 $REDUCE(M_{1})$ and reassign $S$; /* apply $R3$ and $R4$ */
@@
Fast Calculation of Move Gain. For each vertex $u\in V$, we maintain $B_{u}=w(u)-\sum_{v\in S,\{u,v\}\in E_{r}}p(\{u,v\})$. Then, the move gain of ADD(v)ADD(v)—the gain of the net benefit after applying a $ADD(v)$, is $B_{v}$, and the move gain of $SWAP(u,v)$ is $B_{v}-B_{u}$. After an $ADD(v)$ move, for each $u\in N_{r}(v)$, we update $B_{u}$ as $B_{u}\coloneqq B_{u}-p(\{u,v\})$. After a $SWAP(u,v)$ move, for each $x\in N_{r}(u)$, $B_{x}\coloneqq B_{x}+p(\{x,u\})$ and for each $y\in N_{r}(v)$, $B_{y}\coloneqq B_{y}-p(\{y,v\})$. The time complexity of performing $ADD$ or $SWAP$ is $O(\Delta)$, and of REDUCE is $O(|E|)$. The overall complexity of applying all three operators is $O(|E|)$. Notably, the REDUCE operator is not limited to the GIS problem and can enhance the efficiency of local search in other optimization problems.
Neighborhood Search with Tabu and Perturbation. The neighborhood search in the RLS iteratively performs the three basic operators above to improve the current solution $S$, as shown in Algorithm 4.
At each iteration in lines 4-4, it first builds $M_{1}$, the set of vertices that can be added into $S$ by $ADD$, and $M_{2}$, the set of vertices that can be swapped into $S$ by $SWAP$. By evaluating $M_{1}$ and $M_{2}$, $ADD$ or $SWAP$ with the largest move gain is applied, as presented in lines 4-8. After performing $ADD$ or $SWAP$, we additionally use the REDUCE operator to add more vertices to $S$. This narrows the search space $M_{1}=V\setminus N_{p}[S]$ of $ADD$.
To avoid short-term search cycles, a tabu list is used to exclude recently visited vertices for successive iterations (line 4). The tabu tenure for a vertex is set to $10+random(0,|M_{2}|)$, as commonly used in local search algorithms (Wu et al. 2023, Zhou and Hao 2017, Jin and Hao 2015). To further avoid local optima, a perturbation step is applied during neighborhood search (lines 4-4). Specifically, when we detect that no operators can be applied or the best solution remains unchanged after a certain number of iterations, we remove a portion of low-contribution vertices. The whole neighborhood search terminates when the consecutive number of iterations without improvement reaches the search tolerance parameter $L$ (line 4).